# Multiple Choice Learning of Low Rank Adapters for Language Modeling

### v1

@@ 4 Theoretical Analysis
We justify the use of MCL in the context of language modeling by assuming that the sequence distribution can be expressed as a mixture. We contrast this formulation with standard maximum likelihood estimation (1). In particular, Section 4.1 establishes a connection between our approach and the Expectation-Maximization (EM) algorithm and derives lower and upper bounds on the optimal achievable test loss. We then apply this analysis to the case of Markov chains in Section 4.2, where the dynamic of the method is simulated in a controlled setting in Section 4.3.
@@ 4.1 Training dynamics and optimality conditions
For the next-token prediction loss (1), one can show that: $\mathrm{min}_{\theta}\;\mathcal{L}(\theta)=\mathcal{H}(x\,|\,c)$, where $\mathcal{H}(x\,|\,c)\triangleq-\mathbb{E}_{c,x}[\mathrm{log}\;p(x\,|\,c)]$ is the entropy of $p$ mackay2003information . Following the rationale of Section 3.1, let us now assume the data distribution can be written as a mixture. In the case of the WTA loss under this setup, we have the following proposition.
Proposition 1.
Assume the data-generating process is $p(x\,|\,c)=\sum_{k=1}^{K}p(z_{k}\,|\,c)\;p(x\,|\,z_{k},c)$ (Asm. 2) with the number of modes equal to the number of hypotheses. Also assume perfect model expressiveness (Asm. 1), and large enough batch size to approximate the true risk well (Asm. 3). Then:
@@
Proof. See Appendix B.∎
$(i)$ in Proposition 1 describes the relationship between LoRA-MCL and the hard-EM algorithm. $(ii)$ provides an expression for the WTA loss as a conditional entropy, $\mathcal{H}(x\,|\,c,z)$, under the assumption of a perfect matching between the hypotheses and the modes. Proposition 1 also establishes in $(iii)$ both a lower and an upper bound on the optimal achievable loss for LoRA-MCL, given in $(a)$ and $(b)$, respectively. Note that the gap between these bounds is $\mathcal{H}(x\,|\,c,z)-\mathrm{min}_{\theta}\,\mathcal{L}(\theta)+\mathrm{log}\,K=-\mathbb{E}_{c}[\mathcal{I}(x,z\,|\,c)]+\mathrm{log}\,K$ where $\mathcal{I}$ denotes the mutual information. Finally, $(c)$ shows in particular that $\mathrm{min}_{\theta}\,\mathcal{L}^{\mathrm{WTA}}(\theta)\leq\mathrm{min}_{\theta}\,\mathcal{L}(\theta)$.
@@ 4.2 Case of Markov chains
To make the analysis more concrete, we consider the case where the sequence of tokens is generated from Markov chains. Formally, given a finite state space $\mathcal{V}$, a homogeneous Markov chain is defined by an initial distribution $\pi$ over $\mathcal{V}$ and a transition matrix $P\in[0,1]^{\mathcal{V}\times\mathcal{V}}$. A sequence $(x_{1},\dots,x_{T})$ is sampled from the Markov chain if $x_{1}\sim\pi$ and, for each $t\geq 1$, the conditional distribution of $x_{t+1}$ given $x_{t}$ is $x_{t+1}\,|\,x_{t}\sim P_{x_{t},\cdot}$, where $P_{i,j}=\mathbb{P}(x_{t+1}=j\,|\,x_{t}=i)$. In the following, we ignore the initial warm-up phase and we assume that $\pi$ is the stationary distribution of $P$, i.e., such that $\pi P=\pi$. In this case, we will denote $x\sim\mathrm{MC}(P)$.
While the study of the training dynamics of transformers on Markov Chain data has been investigated in previous works edelman2024evolution ; rajaraman2024analysis ; makkuva2024attention ; zekri2024large , our setup instead considers a mixture of Markov chains gupta2016mixtures ; kausik2023learning . Assuming a uniform mixture, the data generating process is $x\sim\frac{1}{K}\sum_{k=1}^{K}\mathrm{MC}(P_{k})$, which is equivalent to:
$z_{k}\sim\mathcal{U}(1,\dots,K),\quad x\,|\,z_{k}\sim\mathrm{MC}(P_{k})\;.$ (9)
@@
Corollary 1.
As per Assumption 5, let us assume that the data-generating process writes as a uniform mixture of Markov chains of order $n=1$. We denote as $\hat{P}(\theta)\triangleq(p(x_{t+1}=j\,|\,x_{t}=i))_{i,j}$ the predicted transition matrix when using a language model with parameters $\theta$. Under the same assumptions that in Proposition 1, we have that:
@@
    The inequality (8) holds in this context, where the conditional entropy $\mathcal{H}(x\,|\,z)$ can be computed by weighting the entropy rate of each Markov Chain:
    $\mathcal{H}(x\,|\,z)=-{\color[rgb]{0,0,0}T}\sum_{k=1}^{K}\sum_{i=1}^{|\mathcal{V}|}(\pi_{k})_{i}\Big{[}\sum_{j=1}^{|\mathcal{V}|}(P_{k})_{i,j}\;\mathrm{log}(P_{k})_{i,j}\Big{]}\,.$ (11)
Proof. See Appendix C. ∎
The entropy $\mathcal{H}(x)$, which is theoretical optimal loss of the MLE baseline, can be computed either exactly for short sequences, or approximated, e.g., through Monte-Carlo integration.
Note that we considered first-order Markov chains in our analysis. We expect the properties generalize to higher orders ($n>1$) by using transition matrices in the form $P\in[0,1]^{\mathcal{V}^{n+1}}$ where $P_{i_{1},\dots,i_{n+1}}=p(x_{t+1}=i_{n+1}\,|\,x_{t}=i_{n},\dots,x_{t-n+1}=i_{1})$. Please refer to Appendix C for further discussions.
@@ 4.3 Illustration with synthetic data
Figure 1: Comparison of LoRA-MCL with vanilla maximum likelihood estimation (MLE). The experimental setup is detailed in Section 4.2. (Left) Validation loss over training steps (averaged across three random seeds) for LoRA-MLE (blue) and LoRA-MCL (red). The theoretical optimal MLE loss, approximated via Monte Carlo sampling, is given by the entropy $\mathcal{H}(x)$. The gray shaded region represents the lower and upper bounds of the theoretical optimal MCL loss, as given by $(a)$ and $(b)$ in (8). (Right) Learned transition matrices (top) are compared with the references (bottom). MLE converges approximately toward the weighted average $\bar{P}$ defined as the right-hand side of (10). In contrast, LoRA-MCL successfully recovers the two underlying modes.
To illustrate our approach, let us evaluate our algorithm on a synthetic dataset, for which results are given in Figure 1. We used $\mathcal{V}=\{1,2\}$, and $(P_{1},P_{2})=(P(p_{1},q_{1}),P(p_{2},q_{2}))$, with $P(p,q)\triangleq\left[\begin{array}[]{cc}1-p&p\\ q&1-q\end{array}\right]\;$ and $p,q\in[0,1]$. We sampled data from a mixture of two Markov chains following (9): we generated a sequence by sampling first $k\sim\mathcal{U}\{1,2\}$ and then $P_{k}$. Once $P_{k}$ was set, we sampled uniformly the initial state, then sampled the Markov chain according to the transition matrix until reaching the maximum sequence length ($T=32$ here).
We considered a GPT-2-like architecture radford2019language ; gpt-neo using local-attention suggested by makkuva2024attention to improve convergence on Markov chain data. We then trained the model with $1$ and $2$ hypotheses, using LoRA adapters as described above, $1$ hypothesis corresponding to vanilla maximum likelihood estimation. Training details are provided in Appendix D. Figures 1 and 4 show both the evolution of the losses along training and the predicted transition matrix from the trained models. We see that the optimum is close to being global in this setup. While LoRA-MCL matches can capture the two modes of the mixture, we see that MLE tends to predict the weighted average of the transition matrices given by (10), which is consistent with Prop. 1.
@@ 5 Empirical evaluation
We evaluate LoRA-MCL on realistic datasets and large-scale models for audio and image captioning tasks. Predicting a textual description for images or audio signals is an ill-posed problem: from an input image or audio clip, multiple descriptions may be plausible. We argue this is a particularly relevant real-world case of ill-posed problems, where the conditional output distribution is inherently multi-modal. Our goal is to demonstrate that LoRA-MCL provides a competitive approach for capturing these distribution modes when fine-tuning either audio (Sec. 5.2) or vision-language models (Sec. 5.3). We present hereafter the experimental setup. Please refer to Appendix E.2 for further details.
5.1 Experimental Setup
@@
Datasets. We experimented on both Clotho-V2 font2013freesound ; drossos2020clotho , and AudioCaps gemmeke2017audio ; kim2019audiocaps datasets for the audio captioning task, while we make use of the TextCaps dataset sidorov2020textcaps for a challenging task for vision language models coined as image captioning with reading comprehension. Table 4 of the Appendix describes the datasets sizes.
@@
Architecture and training details with audio data. We used the instructed version of Qwen2-Audio chu2024qwen2 as the base model, which has ${\sim}8.4$ billion parameters and a vocabulary size $|\mathcal{V}|=156{,}032$. We used LoRA adapters applied to the $Q,K,V$ linear projections of the attention modules, and the upside and downside projections of the feedforward blocks, for all the layers of the transformers. We used a rank $r$ and scaling factor $\alpha$, with $r=\alpha=8$ unless otherwise stated. We trained for 1 epoch on AudioCaps and 10 epochs on Clotho with a validation loss computed at the end of each epoch.
@@
Architecture and training details with visual data. We used LLaVA 1.6 liu2023visual with Vicuna-7B zheng2023judging for the LLM, as the base model which features ${\sim}7.1$ billion parameters and a vocabulary size $|\mathcal{V}|=32{,}000$. We used the official codebase for the implementation. Following standard practices zhou2024empirical , we applied LoRA adapters only for the LLM decoder. The adapters were applied to $Q,K,V$, upside and downside projections as in Qwen2-Audio, and we used $r=\frac{\alpha}{4}=8$ unless otherwise stated. Training was done over 1 epoch (without validation data), and the validation set of TextCaps was used for evaluation.
@@
Metrics. We evaluate the approach using various quality metrics. First, we compute the test-loss (or perplexity), which evaluates the likelihood of the reference captions under the trained model. Second, we use natural language generation quality metrics (BLEU-$n$, ROUGE, METEOR) papineni2002bleu ; lin2004rouge ; banerjee2005meteor based on $n$-gram matching between predicted and target sentences. We also leverage the latter proposed CIDEr, SPICE and SPIDEr vedantam2015cider ; anderson2016spice ; liu2017improved metrics, which show higher correlation with human evaluation for captioning tasks. Finally, we use Sentence-BERT (sBERT) reimers2019sentence ; zhou2022can for embedding comparison. We consider sentence-based oracle evaluation for these metrics (see e.g., (lee2016stochastic, , Sec. 4.3) or labbe2022my ). To evaluate the diversity of the generated captions, we consider: Div-$n$, which compares the number of distinct $n$-grams against the number of $n$-grams for a set of generated captions as in mei2022diverse and Self-BLEU (mBLEU-4), which computes BLEU similarity scores between generated captions zhu2018texygen .
@@
Baselines. To validate our approach, we compare it against a maximum likelihood baseline (LoRA-MLE) trained using (1), under the same conditions as the ones considered for our multi-hypothesis model. Specifically, both models use the same LoRA configuration, the same number of trainable parameters (the LoRA rank for the baseline is $K\times$ larger to this end), and the same number of iterations. At inference time, for each decoding method applied to the baseline that returns $K$ sentences, we decode the same number of candidates with LoRA-MCL. When evaluating Maximum A Posteriori methods –greedy, beam search beamSearch and diverse beam search vijayakumar2018diverse –, we ensure a consistent computational budget by aligning the number of forward passes. For example, if the MLE baseline uses a beam size of $B$, our model uses a beam size of $\frac{B}{K}$ per hypothesis.
@@ 5.2 Audio captioning
Figure 2: Quality vs. Diversity on Audio Captioning with $5$ candidates. Quality measured by SPIDEr ()($\uparrow$)( ↑ ) and Diversity by mBLEU-4 ()($\downarrow$)( ↓ ). Marker shape stands for the method, and size is proportional to the number of forwards performed per example at inference time. Note that LoRA-MLE was trained with two rank values, $r=8$ and $r=8K$, for fair comparison in terms of parameter count. Color corresponds to the decoding method (beam search or diverse beam search). Values of $\varepsilon$ and $\lambda$ for LoRA-MCL and diverse beam search, resp., are indicated in the plot, with the color shade proportional to the corresponding parameter value to better distinguish the markers.
Quality vs. diversity trade-off. Quality–diversity performance is shown in Figure 2. While a well-chosen value of $\lambda$ allows Diverse Beam Search applied to LoRA-MLE to be competitive, LoRA-MCL (circles) achieves the best trade-off between quality and diversity, appearing in the top-left corner of the plot. Although increasing the beam size generally improves performance for standard beam search, we observe that increasing the beam size within each group negatively impacts Diverse Beam Search. In theory, setting $\varepsilon=0$ should maximize both quality and diversity. However, using a small positive value of $\varepsilon$ ensures all heads are trained, albeit at the cost of reduced diversity and potentially quality when $\varepsilon$ is too large. Full results are given in Tables 5 and 6 of the Appendix.
Effect of the number of hypotheses. We include in Figures 6 and 5 (given in the appendix) the same plot as in Figure 2, but using 3 and 4 candidates, respectively. These results further reinforce the findings obtained with 5 hypotheses. In Table 5.2, we report the evolution of the test negative log-likelihood (NLL) loss as a function of the number of hypotheses, showing a monotonically decreasing trend. This provides additional evidence for Proposition 1, and shows that LoRA-MCL improves its coverage of the data distribution modes as $K$ increases.
Table 1: Test Loss ()($\downarrow$)( ↓ ) as a function of the number of hypotheses $K$. LoRA-MCL is trained with $\varepsilon=0.05$ and $r=8$.
@@
LoRA-MLE ($r=8$) 1 2.152 2.811
LoRA-MLE ($r=40$) 1 2.124 2.902
LoRA-MCL 2 2.071 2.700
LoRA-MCL 3 2.021 2.665
LoRA-MCL 5 1.927 2.629
LoRA-MCL 8 1.869 2.592
Table 2: Quality and Diversity Evaluation on TextCaps with $3$ candidates. Best scores are in bold, second-best are underlined. For each of the presented metrics, higher is better ()($\uparrow$)( ↑ ) except for mBLEU-4 ()($\downarrow$)( ↓ ). LoRA-MCL is trained with $\varepsilon=0.1$, $r=8$ and $\alpha=32$. LoRA-MLE is trained with $r=24$ and $\alpha=4\times r=96$, ensuring the same dynamics and number of parameters across models. For completeness, we also trained LoRA-MLE with $r=8$ and $\alpha=32$ in the rows marked with .
Training Decoding Beam mBLEU-4 BLEU-4 METEOR sBERT CIDEr-D SPICE SPIDEr
LoRA-MLE Beam Search 3 0.688 0.318 0.315 0.670 1.517 0.244 0.873
LoRA-MLE Beam Search 6 0.786 0.338 0.326 0.671 1.557 0.246 0.895
LoRA-MLE DBS ($\lambda=0.8$) 3 0.437 0.349 0.327 0.686 1.590 0.251 0.909
LoRA-MLE DBS ($\lambda=1.0$) 3 0.416 0.348 0.326 0.685 1.586 0.250 0.906
LoRA-MLE DBS ($\lambda=1.0$) 3 0.425 0.357 0.327 0.686 1.601 0.252 0.915
LoRA-MLE DBS ($\lambda=0.8$) 6 0.671 0.341 0.328 0.681 1.573 0.251 0.903
LoRA-MLE DBS ($\lambda=1.0$) 6 0.666 0.340 0.328 0.680 1.577 0.250 0.904
LoRA-MLE DBS ($\lambda=1.0$) 6 0.671 0.341 0.328 0.680 1.584 0.251 0.908
LoRA-MCL Beam Search 1 0.520 0.344 0.330 0.690 1.674 0.255 0.955
LoRA-MCL Beam Search 2 0.490 0.360 0.333 0.687 1.627 0.258 0.932

### v2

@@ 4 Theoretical Analysis
We justify the use of MCL for language modeling by assuming the sequence distribution is a mixture. Section 4.1 links our method to EM and derives lower and upper bounds on the optimal achievable test loss. We then apply this analysis to the case of Markov chains and simulate the method’s dynamics in a controlled setting.
@@ 4.1 Training dynamics and optimality conditions
For the next-token prediction loss in (1), one can show that: $\mathrm{min}_{\theta}\;\mathcal{L}(\theta)=\mathcal{H}(x\,|\,c)$, where $\mathcal{H}(x\,|\,c)\triangleq-\mathbb{E}_{c,x}[\mathrm{log}\;p(x\,|\,c)]$ is the entropy of $p$ (MacKay, 2003). Following the rationale of Sec. 3.1, let us now assume the data distribution can be written as a mixture. In the case of the WTA loss, we have the following proposition.
Proposition 1 (Proof in Apx. B).
Assume a data-generating process $p(x\,|\,c)=\sum_{k=1}^{K}p(z_{k}\,|\,c)\;p(x\,|\,z_{k},c)$ (Asm. 2), perfect model expressiveness (Asm. 1), and large enough batch size to approximate the true risk (Asm. 3). Then:
@@
    Assuming disjoint components for the data mixture (Asm. 4), and assuming $p(x\,|\,z_{k},c)=p(x\,|\,c;\theta_{k})$ for each $k$, then
    $\mathcal{L}^{\mathrm{WTA}}(\theta)=\mathcal{H}(x\,|\,c,z)\,$, (6)
    where $\mathcal{H}(x\,|\,c,z)\triangleq\mathbb{E}_{c}\left[\sum_{k=1}^{K}p(z_{k}\,|\,c)\mathcal{H}\big(x\,|\,c,z_{k}\big)\right]$ is the conditional entropy given the random variable $z$.
@@
    $\begin{split}\min_{\theta}\,\mathcal{L}(\theta)-\log K\overset{(a)}{\leq}\min_{\theta}\,\mathcal{L}^{\mathrm{WTA}}(\theta)\\ \overset{(b)}{\leq}\mathcal{H}(x\,|\,c,z)\overset{(c)}{\leq}\min_{\theta}\,\mathcal{L}(\theta)\;.\end{split}$ (7)
@@
$(i)$ in Prop. 1 describes the relationship between LoRA-MCL and the hard-EM algorithm. $(ii)$ provides an expression for the WTA loss as a conditional entropy, $\mathcal{H}(x\,|\,c,z)$, assuming a perfect matching between the hypotheses and the modes. $(iii)$ establishes both a lower and an upper bound on the optimal achievable loss for LoRA-MCL, given in $(a)$ and $(b)$, respectively. Note that the gap between these bounds is $\mathcal{H}(x\,|\,c,z)-\mathrm{min}_{\theta}\,\mathcal{L}(\theta)+\mathrm{log}\,K=-\mathbb{E}_{c}[\mathcal{I}(x,z\,|\,c)]+\mathrm{log}\,K$ where $\mathcal{I}$ denotes the mutual information. Finally, $(c)$ shows in particular that $\mathrm{min}_{\theta}\,\mathcal{L}^{\mathrm{WTA}}(\theta)\leq\mathrm{min}_{\theta}\,\mathcal{L}(\theta)$.
@@ 4.2 Case of Markov chains
To make the analysis more concrete, we consider the case where the sequence of tokens is generated from Markov chains. Given a finite state space $\mathcal{V}$, a homogeneous Markov chain is defined by an initial distribution $\pi$ over $\mathcal{V}$ and a transition matrix $P\in[0,1]^{\mathcal{V}\times\mathcal{V}}$. A sequence $(x_{1},\dots,x_{T})$ is sampled from the Markov chain if $x_{1}\sim\pi$ and, for each $t\geq 1$, $x_{t+1}\,|\,x_{t}\sim P_{x_{t},\cdot}$, where $P_{i,j}=p(x_{t+1}=j\,|\,x_{t}=i)$. In the following, we ignore the initial warm-up phase, and we assume that $\pi$ is the stationary distribution of $P$. In this case, we will denote $x\sim\mathrm{MC}(P)$.
While the study of the training dynamics of transformers on Markov Chain data has been investigated in previous works (Edelman et al., 2024; Rajaraman et al., 2024; Makkuva et al., 2024; Zekri et al., 2024), our setup instead considers a mixture of Markov chains (Gupta et al., 2016; Kausik et al., 2023). Assuming a uniform mixture, the data generating process is $x\sim\frac{1}{K}\sum_{k=1}^{K}\mathrm{MC}(P_{k})$;
$k\sim\mathcal{U}(1,\dots,K),\quad x\,|\,z_{k}\sim\mathrm{MC}(P_{k})\;.$ (8)
@@
Corollary 1 (Proof in Apx. C).
Assume the data-generating process is a uniform mixture of first-order Markov chains. Let $\hat{P}(\theta)\triangleq(p_{\theta}(x_{t+1}=j\,|\,x_{t}=i))_{i,j}$ be the predicted transition matrix. Under the same Asm. as in Prop. 1:
@@
    When the MLE estimator trained with (1) reaches its optimal loss, we have
@@
    The inequalities (7) holds, where $\mathcal{H}(x\,|\,z)$ is a weighted average of the entropy rate of each Markov Chain.
The entropy $\mathcal{H}(x)$, which is the optimal loss of the MLE baseline, can be computed either exactly for short sequences or approximated, e.g., through Monte-Carlo integration. Our analysis considers first-order Markov chains, but we expect the results to extend to higher orders (see Apx. C).
@@ 4.3 Illustration with synthetic data
Figure 2: Comparison of MCL with MLE. (Left) Validation loss over training steps (averaged across three seeds) for LoRA-MLE (blue) and LoRA-MCL (red). The theoretical optimal MLE loss is the entropy $\mathcal{H}(x)$. The gray shaded region represents the bounds of the theoretical optimal MCL loss, as given by $(a)$ and $(b)$ in (7). (Right) Learned transition matrices (top) versus references (bottom). MLE converges approximately toward the weighted average $\bar{P}$ (right-hand side of (9)), whereas LoRA-MCL recovers the two modes.
To illustrate our approach, let us evaluate our algorithm on a synthetic dataset, for which results are given in Fig. 2. We used $\mathcal{V}=\{1,2\}$, and $(P_{1},P_{2})=(P(p_{1},q_{1}),P(p_{2},q_{2}))$, with $P(p,q)\triangleq\left[\begin{array}[]{cc}1-p&p\\ q&1-q\end{array}\right]\;$ and $p,q\in[0,1]$. We sampled data from a mixture of two Markov chains following (8): we generated a sequence by sampling first $P_{k}$ with $k\sim\mathcal{U}\{1,2\}$. Once $P_{k}$ was set, we sampled the initial state uniformly, then we sampled the Markov chain according to the transition matrix until reaching the maximum sequence length ($T=32$ here).
We considered a GPT-2-like architecture (Radford et al., 2019; Black et al., 2021) using local-attention suggested by Makkuva et al. (2024) to improve convergence on Markov chain data. We then trained the model with one and two hypotheses, using LoRA adapters as described above, one hypothesis corresponding to vanilla MLE. Training details are provided in Apx. D. Figures 2 and 7 show both the evolution of the losses along training and the predicted transition matrix from the trained models. We see that the optimum is close to being global in this setup. While LoRA-MCL matches can capture the two modes of the mixture, we see that MLE tends to predict the weighted average of the transition matrices given by (9), which is consistent with Prop. 1.
@@ 5 Empirical evaluation
We evaluate LoRA-MCL on realistic datasets and large-scale models for audio and image captioning tasks, as well as machine translation. Predicting a textual description for images or audio signals is an ill-posed problem: from an input image or audio clip, multiple descriptions may be plausible; this is a real-world case where the conditional output distribution is inherently multi-modal. Similarly, Machine Translation is a one-to-many problem (Ott et al., 2018). We demonstrate that LoRA-MCL provides a competitive approach for capturing these distribution when fine-tuning either audio, vision-language, or language-only models. We describe the experimental setup. See Apx. E.6 for details.
5.1 Experimental Setup in captioning
Figure 3: Quality–diversity trade-off on audio captioning (5 candidates). SPIDEr $(\uparrow)$ for quality, mBLEU-4 $(\downarrow)$ for diversity. Marker shape stands for the method, color for the decoding method, and size is proportional to forward passes per example at inference. LoRA-MLE uses $r\in\{8,8K\}$ for parameter parity. LoRA-MCL uses circle markers: Relaxed (black edge) and Annealed (red edge).
@@ 5.1 Experimental Setup in captioning
Datasets. We experimented on both Clotho-V2 (Font et al., 2013; Drossos et al., 2020), and AudioCaps (Gemmeke et al., 2017; Kim et al., 2019) datasets for the audio captioning task, while we make use of the TextCaps dataset (Sidorov et al., 2020) for the task of image captioning with reading comprehension. Table 4 describes the datasets sizes.
@@
Experimental details in audio. We used the instructed Qwen2-Audio (Chu et al., 2024) as the base model, which has ${\sim}8.4$ billion parameters and a vocabulary size $|\mathcal{V}|=156{,}032$. We used LoRA adapters applied to the $Q,K,V$ linear projections of the attention modules, and the upside and downside projections of the feedforward blocks, across all layers. We used a rank $r$ and scaling factor $\alpha$, with $r=\alpha=8$ unless otherwise stated. We trained for 1 and 10 epochs on AudioCaps and Clotho, respectively.
@@
Experimental details with visual data. We used LLaVA 1.6 (Liu et al., 2023), as the base model which features ${\sim}7.1$ billion parameters and a vocabulary size $|\mathcal{V}|=32{,}000$. We applied LoRA adapters only for the LLM decoder following (Zhou et al., 2024). The adapters were applied to $Q,K,V$, upside and downside projections as in Qwen2-Audio, and we used $r=\frac{\alpha}{4}=8$ unless otherwise stated. Training was done over 1 epoch (without validation data), and the validation set of TextCaps was used for evaluation.
@@
Metrics. We evaluate both quality and diversity. For quality, we report test-loss and standard Natural Language Generation metrics (BLEU, ROUGE, METEOR) (Papineni et al., 2002; Lin, 2004; Banerjee and Lavie, 2005), and captioning-specific scores (CIDEr, SPICE, and SPIDEr) (Vedantam et al., 2015; Anderson et al., 2016; Liu et al., 2017), which better correlate with human judgments in captioning. We also use Sentence-BERT (Reimers, 2019). We consider sentence-based oracle evaluation for these metrics (see Lee et al. (2016); Labbé et al. (2022)). For diversity, we used mBLEU-4 (Mei et al., 2022b) measuring similarity across generated captions (Zhu et al., 2018).
@@
Baselines. We compare LoRA-MCL against the MLE baseline (LoRA-MLE) trained using (1), under the same conditions as the ones considered for our multi-hypothesis model. Specifically, both models use the same LoRA configuration, the same number of trainable parameters (the LoRA rank for the baseline is $K\times$ larger to this end), and the same number of iterations. We also considered a Mixture of Low Rank Experts (Muqeeth et al., 2024; Wu et al., 2024; Li et al., 2024) (LoRA-MoE) as a baseline. See Apx. E.3 for more details on the training methods. At inference time, for each decoding method applied to the baseline that returns $K$ sentences, we decode the same number of candidates with LoRA-MCL. When evaluating MAP methods such as greedy, beam search (BS) (Lowerre, 1976), and diverse beam search (DBS) (Vijayakumar et al., 2018), we ensure a consistent computational budget by aligning the number of forward passes. In this case, if LoRA-MLE or LoRA-MoE uses a beam size of $B$, our model uses a beam size of $\frac{B}{K}$ per hypothesis. Finally, we experimented Test-Time Augmentation (TTA) (Wang et al., 2019) with LoRA-MLE in Audio Captioning, applying SpecAugment (Park et al., 2019) $K$ times to the input Spectrogram to expect diverse outputs.
@@ 5.2 Audio captioning
Quality vs. diversity trade-off. Quality–diversity performance is shown in Fig. 3. For readability, only the TTA runs (orange diamonds) with optimal augmentation strength on the quality–diversity are displayed in Fig. 3. It is strong on Clotho, but performs poorly on AudioCaps. We notice that a well-chosen value of $\lambda$ allows DBS applied to LoRA-MLE and LoRA-MoE to be competitive. LoRA-MCL (circles) achieves the best trade-off between quality and diversity, appearing in the top-left corner of the plot, where the best relaxation technique (annealed or relaxed) depends on the setup. Although increasing the beam size generally improves performance for standard beam search, we observe that increasing the beam size within each group in DBS can negatively impact DBS, as observed in Clotho. Additional results and comparisons are provided Apx. E.6.4.
Table 1: Test Loss $(\downarrow)$ as a function of $K$. LoRA-MCL is trained using $\varepsilon=0.05$ and $r=8$.
@@
LoRA-MLE ($r=8$) 1 2.203 2.812
LoRA-MLE ($r=8\times 3$) 1 2.195 2.868
LoRA-MLE ($r=8\times 7$) 1 2.182 2.935
LoRA-MCL 3 2.063 2.663
LoRA-MCL 5 1.999 2.643
LoRA-MCL 7 1.932 2.612
Effect of the number of hypotheses. Table 1 reports the test negative log-likelihood as a function of $K$. The monotonically decreasing trend provides further evidence for Prop. 1, indicating that LoRA-MCL achieves better coverage of the data distribution modes as $K$ increases.